# Second- and third-order properties of multidimensional Langevin equations

### v9

@@ 7.6. Second-order covariance functions to quadratic order
where the base case $n=1$ is assumed for Eq. (277), and for Eq. (278) follows from our assumptions that $\mathbf{D}^{+-}$ and $\langle\mathbf{x}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle$ are $a^{2}\mathcal{O}((a,b)^{2})$, and that $\langle\mathbf{x}^{(2)\pm}{\mathbf{x}^{\mp}}^{\mathsf{T}}\rangle=a\mathcal{O}((a,b)^{2})$. Using the properties stated in the subsection “Nonlinear drift”, the induction step is readily performed. For the case of an integrated variable $y$ of even parity, we additionally need the equalities in the subsection “Integrated variables” and:
@@
We also need to use Eq. (276) and:
(280) ${{}\widehat{\mathbf{A}}}^{(2)}_{(2)}\langle\mathbf{x}^{(2)}\rangle+{{}\widehat{\mathbf{A}}}^{(2)}_{(3)}\langle\mathbf{x}^{(3)}\rangle=a^{2}\mathcal{O}((a,b)^{2}).$
@@ 8.1. Linear Gaussian processes: Covariance functions and detailed balance
(282) $\langle\boldsymbol{\xi}(t){\boldsymbol{\xi}(t^{\prime})}^{\mathsf{T}}\rangle$ $=2\mathbf{D}\delta(t-t^{\prime}).$
The first property of interest is:
@@
By similar reasoning, the second derivative of the position-position covariance function is (negative) the velocity-velocity covariance function:
@@ 8.2. Almost Markovian dynamics
Following the discussion in the first section (“Linear Gaussian stationary systems”), we may judge the significance of this quantity by comparing to $2\mu^{-1}\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{D}{\mathbf{A}_{\mathbf{v}}^{-1}}^{\mathsf{T}}$. The last covariance of interest is:
@@
(308) $\langle\mathbf{v}(\tau)\mathbf{v}(0)^{\mathsf{T}}\rangle=\exp(\mu\mathbf{A}_{\mathbf{v}}\tau)\langle\mathbf{v}\mathbf{v}^{\mathsf{T}}\rangle+\mathcal{O}(\mu^{-2}),$
but as a consequence of the twice differentiability of the position-position covariance function, it must satisfy:
(309) $\int_{0}^{\infty}\mathrm{d}\tau\,\langle\mathbf{v}(\tau)\mathbf{v}(0)^{\mathsf{T}}\rangle+\int_{0}^{\infty}\mathrm{d}\tau\,\langle\mathbf{v}(-\tau)\mathbf{v}(0)^{\mathsf{T}}\rangle=\mathbf{0},$
if $\mathbf{x}$ is stationary, meaning that the $\mathcal{O}(\mu^{-2})$ correction has an integral of value $\mathcal{O}(\mu^{-1})$.
@@ 8.3. Third-order properties
$\langle x^{i}v^{j}v^{k}\rangle$, $\widetilde{L}(x^{i},v^{j},v^{k})$, $\langle v^{i}v^{j}v^{k}\rangle$, $L(v^{i}v^{j},v^{k})$ (satisfying Eq. (85)), $\langle x^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$, and $\langle v^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$. With the exception of the last two, these quantities can be straightforwardly estimated from time-lapse data without corrections due to fluctuations. The estimation of fluctuation covariance $\mathcal{B}$ and measurement error covariance $\boldsymbol{\Lambda}$ is derived in the Supplementary Material in [10]. However, Eq. (S70) should read $(6/11)(-1,1,1,-3,1,1)^{\mathsf{T}}$ as in the code. We may write these estimators in terms of the measured values $\mathbf{y}:=\mathbf{x}+\boldsymbol{\eta}$, where $\boldsymbol{\eta}$ is measurement noise. Using the second differences:
@@
where Eq. (339) is substituted into Eq. (338). The resulting equality can be substituted into the expression for $L(x^{i}x^{j},x^{k})$ and the ensemble covariance of angular momenta applied. It should be mentioned that the $\mu^{-2}L(\widetilde{v}^{i}\widetilde{v}^{j},\widetilde{x}^{k})$ and $\mu^{-3}L(\widetilde{v}^{i}\widetilde{v}^{j},\widetilde{v}^{k})$ terms contribute only $\mathcal{O}(\mu^{-4})$ to the ensemble covariance of $L(x^{i}x^{j},x^{k})$. For $\widetilde{L}(x^{i},v^{j},v^{k})$ and $L(v^{i}v^{j},v^{k})$, we can use the usual prescription without issue. It is worth noting that using the above equality and using the prescription for the angular momenta $L(\widetilde{x}^{i}\widetilde{v}^{j},\widetilde{x}^{k})$ gives a different expression for the ensemble covariance:
@@
For the inhomogeneous diffusion, we have
@@ 9. Conclusions
We have used Koopman eigenfunctions to facilitate calculations in the case of integrated variables (written in terms of martingales) and nonlinear drift. In the latter case, we might consider these to be new coordinates under a nonlinear coordinate transformation, as these quantities obey a law of linear drift. However, the time-reversed dynamics of Koopman eigenfunctions do not obey a law of linear drift, and hence our use of them is restricted to being a calculational tool.
We have introduced criteria for judging quantitative significance, both for interesting effects such as non-Gaussianity or time-irreversibility, and for comparing experiment with theory. Our analysis takes linear Gaussian models as a reference point and utilizes second- and third-order covariance functions. Due to moment closure problem, our analysis is limited to asymptotic expansion in the drift nonlinearity. For strongly nonlinear drift, other methods may be more suitable [23].
@@
We expect that the framework presented in this work will be useful in analyses of high-dimensional stochastic systems. We note that extensions to our analysis are possible only up to fourth order (cubic polynomial for drift and quadratic polynomial for diffusion). Afterwards, the analysis procedure may fail (see Appendix B). Besides, analysis of moments beyond fourth order may be infeasible in biological applications. Lastly, future directions would extend this approach to include temporal [25] and population [26] heterogeneity.

### v11

@@ 7.6. Second-order covariance functions to quadratic order
where the base case $n=1$ is assumed for Eq. (291), and for Eq. (292) follows from our assumptions that $\mathbf{D}^{+-}$ and $\langle\mathbf{x}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle$ are $a^{2}\mathcal{O}((a,b)^{2})$, and that $\langle\mathbf{x}^{(2)\pm}{\mathbf{x}^{\mp}}^{\mathsf{T}}\rangle=a\mathcal{O}((a,b)^{2})$. Using the properties stated in section 7.4, the induction step is readily performed. For the case of an integrated variable $y$ of even parity, we additionally need the equalities in section 7.5 and:
@@
We also need to use, from stationarity:
(294) $0=\frac{\mathrm{d}}{\mathrm{d}\tau}\langle\mathbf{x}^{(2)}\rangle={{}\widehat{\mathbf{A}}}^{(2)}_{(2)}\langle\mathbf{x}^{(2)}\rangle+{{}\widehat{\mathbf{A}}}^{(2)}_{(3)}\langle\mathbf{x}^{(3)}\rangle+a^{2}\mathcal{O}((a,b)^{2}).$
@@ 8.1. Linear Gaussian processes: Covariance functions and detailed balance
(296) $\langle\boldsymbol{\xi}(t){\boldsymbol{\xi}(t^{\prime})}^{\mathsf{T}}\rangle$ $=2\mathbf{D}\delta(t-t^{\prime}),$
where $\boldsymbol{\xi}$ is Gaussian random noise and all coefficients are constant. The first property of interest is:
@@
By similar reasoning, the second derivative of the position-position covariance function is negative the velocity-velocity covariance function:
@@ 8.2. Almost Markovian dynamics
Following the discussion in the first section (“Linear Gaussian stationary systems”), we may judge the significance of this quantity by comparing to $2\mu^{-1}\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{D}{\mathbf{A}_{\mathbf{v}}^{-1}}^{\mathsf{T}}$, i.e., for the “ensemble covariance”:
(319) $\langle x^{i}v^{j}\rangle\langle x^{k}v^{l}\rangle\sim 2\mu^{-2}(\mathbf{A}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}(\mathbf{A}_{\mathbf{v}}^{-1})^{j}_{j^{\prime}}(\mathbf{A}_{\mathbf{v}}^{-1})^{k}_{k^{\prime}}(\mathbf{A}_{\mathbf{v}}^{-1})^{l}_{l^{\prime}}(D^{i^{\prime}k^{\prime}}D^{j^{\prime}l^{\prime}}-D^{i^{\prime}l^{\prime}}D^{j^{\prime}k^{\prime}}),$
where we used, in accordance with Eq. (43):
(320) $L(\widetilde{x}^{i},\widetilde{x}^{j})L(\widetilde{v}^{k},\widetilde{v}^{l})\sim 2(\mathbf{A}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}(\mathbf{A}_{\mathbf{v}}^{-1})^{j}_{j^{\prime}}(D^{i^{\prime}k}D^{j^{\prime}l}-D^{i^{\prime}l}D^{j^{\prime}k}).$
The last covariance of interest is:
@@
(324) $\langle\mathbf{v}(\tau)\mathbf{v}(0)^{\mathsf{T}}\rangle=\exp(\mu\mathbf{A}_{\mathbf{v}}\tau)\langle\mathbf{v}\mathbf{v}^{\mathsf{T}}\rangle+\mathcal{O}(\mu^{-2}).$
However, we also have:
(325) $\int_{0}^{\infty}\mathrm{d}\tau\,\langle v^{i}(\tau)v^{j}(0)\rangle+\int_{0}^{\infty}\mathrm{d}\tau\,\langle v^{i}(-\tau)v^{j}(0)\rangle=0,$
if either $x^{i}$ or $x^{j}$ is stationary, meaning that the $\mathcal{O}(\mu^{-2})$ correction to the velocity-velocity covariance function has an integral of order $\mathcal{O}(\mu^{-1})$.
@@ 8.3. Third-order properties
$\langle x^{i}v^{j}v^{k}\rangle$, $\widetilde{L}(x^{i},v^{j},v^{k})$, $\langle v^{i}v^{j}v^{k}\rangle$, $L(v^{i}v^{j},v^{k})$ (satisfying Eq. (92)), $\langle x^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$, and $\langle v^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$. With the exception of the last two, these quantities can be straightforwardly estimated from time-lapse data without corrections due to fluctuations. The estimation of fluctuation covariance $\mathcal{B}$ and measurement error covariance $\boldsymbol{\Lambda}$ is derived in the Supplementary Material in [10]. However, Eq. (S70) in [10] should read $(6/11)(-1,1,1,-3,1,1)^{\mathsf{T}}$ as in their code. We may write these estimators in terms of the measured values $\mathbf{y}:=\mathbf{x}+\boldsymbol{\eta}$, where $\boldsymbol{\eta}$ is measurement noise. Using the second differences:
@@
where Eq. (355) is substituted into Eq. (354). The resulting equality can be substituted into the expression for $L(x^{i}x^{j},x^{k})$ and the ensemble covariance of angular momenta applied. It should be mentioned that the $\mu^{-2}L(\widetilde{v}^{i}\widetilde{v}^{j},\widetilde{x}^{k})$ and $\mu^{-3}L(\widetilde{v}^{i}\widetilde{v}^{j},\widetilde{v}^{k})$ terms contribute only $\mathcal{O}(\mu^{-4})$ to the ensemble covariance of $L(x^{i}x^{j},x^{k})$. For $\widetilde{L}(x^{i},v^{j},v^{k})$ and $L(v^{i}v^{j},v^{k})$, we can use the usual prescriptions without issue. It is worth noting that using the above equality and using the prescription for the angular momenta $L(\widetilde{x}^{i}\widetilde{v}^{j},\widetilde{x}^{k})$ gives a different expression for the “ensemble covariance”:
@@
For the inhomogeneous diffusion, we have for the “ensemble covariance”:
@@ 8.4. Dynamics on long time-scales
9. Detection of non-Markovianity
Stochastic processes exhibiting temporal heterogeneity have been detected in animal locomotion [24] and cell migration [25]. These may take the form of temporal trends [24] or stochastic variability in the parameters characterizing the stochastic process [25]. The latter case is known as “superstatistics” and usually incorporates variability in the process variance leading to leptokurtic distributions [26, 27]. However, the concept of superstatistics need not involve non-Gaussianity. For example, consider a linear Gaussian process with a fluctuating unobserved mean which itself varies according to a linear Gaussian process. This type of superstatistical process has been used to model cell migration in [28]. In this study, the additional term due to the unobserved process was interpreted as a contribution to the effective noise, leading to a colored noise term.
The most general zero-mean linear Gaussian process in one variable ($x$) corresponding to a linear Markov Gaussian process in two variables, one of which is unobserved ($y$), is given by [29, 30]:
(390) $\begin{pmatrix}\dot{x}\\ \dot{y}\end{pmatrix}$ $=\begin{pmatrix}-\lambda&1\\ -\kappa&0\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}+\boldsymbol{\xi},$
(391) $\langle\boldsymbol{\xi}(t)\rangle$ $=\mathbf{0},$
(392) $\langle\boldsymbol{\xi}(t)\boldsymbol{\xi}^{\mathsf{T}}(t^{\prime})\rangle$ $=2\begin{pmatrix}D&0\\ 0&D^{\prime}\end{pmatrix}\delta(t-t^{\prime}),$
where $\boldsymbol{\xi}$ is Gaussian random noise and all coefficients are constant. For stationarity, we require $\lambda>0$, $\kappa>0$. (We discuss the non-stationary case later.) This formulation allows for feedback $x\to y$, which means that the state of the system could potentially influence the unobserved process. If no such feedback is allowed, then the dynamics instead satisfies:
(393) $\begin{pmatrix}\dot{x}\\ \dot{y}\end{pmatrix}=\begin{pmatrix}-\lambda&1\\ 0&-\eta\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}+\boldsymbol{\xi},$
for constants $\lambda$, $\eta$, which means that the eigenvalues of the dynamical matrix (called $\mathbf{A}$ in previous sections) are constrained to be real. We will hereafter work with the first formulation as the second is a special case.
The Gaussian process $x(t)$ is completely characterized by its covariance function $C(\tau):=\langle x(0)x(\tau)\rangle$:
(394) $C(\tau)=\frac{1}{\mu_{+}-\mu_{-}}\left[{\left({D+\frac{D^{\prime}}{\kappa}}\right)\left({\frac{\mu_{+}}{\lambda}e^{-\mu_{+}\tau}-\frac{\mu_{-}}{\lambda}e^{-\mu_{-}\tau}}\right)-\frac{D^{\prime}}{\kappa}(e^{-\mu_{+}\tau}-e^{-\mu_{-}\tau})}\right]$
($\tau\geq 0$), where $-\mu_{\pm}$ are the eigenvalues of $\mathbf{A}$:
(395) $\mu_{\pm}=\frac{\lambda\pm\sqrt{\lambda^{2}-4\kappa}}{2}.$
We define the correlation function $R(\tau):=C(\tau)/C(0)$. We now compare $R(\tau)$ with the correlation function predicted from a linear Markov process, i.e., $e^{\dot{R}(0)\tau}$. If these two are equal, then the process $x(t)$ is Markovian.
We see that $C(\tau)$ reduces to a single exponential decay when:
(396) $\frac{D}{D^{\prime}}=\frac{1}{\kappa}\left({\frac{\lambda}{\mu_{\pm}}-1}\right).$
This case only occurs for real $\mu_{\pm}$. Writing $\delta_{\mp}$ for the r.h.s. of the above, it can be easily verified that when $\delta_{-}<D/D^{\prime}<\delta_{+}$, we have $\ddot{R}(0)>(\dot{R}(0))^{2}$. Otherwise, if $0\leq D/D^{\prime}<\delta_{-}$ or $\delta_{+}<D/D^{\prime}\leq\infty$, or if $\mu_{\pm}$ are complex conjugates, then we have $\ddot{R}(0)<(\dot{R}(0))^{2}$. Thus, we see that $\ddot{R}(0)$ provides a proxy for non-Markovianity. Although $C(\tau)$ is characterized by four parameters while a single exponential decay is characterized by two parameters, we only need a single equality to test for non-Markovianity. This occurs because when the coefficient of one of the exponential decays is zero, the value of the corresponding decay rate is irrelevant.
Alternatively, instead of $\ddot{R}(0)$, we may consider the integral $\int_{0}^{T}\mathrm{d}\tau\,R(\tau)$ for some $T>0$. As discussed in section 3.3, for the estimation of such integrals from experimental data, $T$ must be significantly less than the trajectory duration. We write:
(397) $R(\tau)=c_{+}e^{-\mu_{+}\tau}+c_{-}e^{-\mu_{-}\tau}$
where $c_{+}+c_{-}=1$. First, we consider the case of $\mu_{\pm}$ real and distinct. If $\delta_{-}<D/D^{\prime}<\delta_{+}$, then $c_{+}>0$ and $c_{-}>0$. Hence, we can use Jensen’s inequality to conclude that $R(\tau)>e^{\dot{R}(0)\tau}$ for $\tau>0$. If $0\leq D/D^{\prime}<\delta_{-}$, then $c_{+}<0$ and $c_{-}>1$, and the inequality $R(\tau)<e^{\dot{R}(0)\tau}$ is equivalent to:
(398) $c_{-}(e^{\Delta\mu\tau}-1)<e^{c_{-}\Delta\mu\tau}-1$
where $\Delta\mu:=\mu_{+}-\mu_{-}$. The above inequality follows immediately from taking derivatives. Similarly, if $\delta_{+}<D/D^{\prime}\leq\infty$, then $c_{+}>1$ and $c_{-}<0$, and the same inequality holds. For the case $\mu_{+}=\mu_{-}=\lambda/2$, we have:
(399) $R(\tau)=(1+c^{\prime}\tau)e^{-\lambda\tau/2},$
where $c^{\prime}$ is a constant with $c^{\prime}\neq 0$ unless $D/D^{\prime}=1/\kappa$. The inequality $R(\tau)<e^{\dot{R}(0)\tau}$ is then equivalent to:
(400) $1+c^{\prime}\tau<e^{c^{\prime}\tau},$
which again follows immediately from taking derivatives.
For the case of $\mu_{\pm}$ complex, we no longer have $R(\tau)<e^{\dot{R}(0)\tau}$ for all $\tau>0$. However, we have:
(401) $\int_{0}^{T}\mathrm{d}\tau\,R(\tau)<\int_{0}^{T}\mathrm{d}\tau\,e^{\dot{R}(0)\tau}$
for all $0<T\leq\infty$. To show this, we note that $R(\tau)$ is the product of a sinusoid and an exponential decay, which implies that:
(402) $\int_{0}^{T}\mathrm{d}\tau\,R(\tau)<\int_{0}^{\tau_{0}}\mathrm{d}\tau\,R(\tau)$
for $\tau_{0}<T\leq\infty$, where $\tau_{0}$ is the first zero of $R(\tau)$. Thus it suffices to show that $R(\tau)<e^{\dot{R}(0)\tau}$ for $0<\tau<\tau_{0}$. Writing $\mu_{\pm}=\lambda/2\pm i\omega$ and $c_{\pm}=(1\pm ic)/2$, this inequality is equivalent to:
(403) $\cos(\omega\tau)+c\sin(\omega\tau)<e^{c\omega\tau},$
which is true because the second derivative of the l.h.s. is negative for $0<\tau<\tau_{0}$, whereas the second derivative of the r.h.s. is non-negative.
Now, we discuss the case where $x(t)$ is integrated of order 1, i.e., $x(0)$ does not possess a stationary distribution but its differences $x(\tau)-x(0)$ do. We also assume that $x(t)$ has been detrended so that $\langle\dot{x}(t)\rangle=0$. We treat this case by setting one of the eigenvalues of $\mathbf{A}$ to zero. We do not set both eigenvalues to zero simultaneously, since this would result in $x(t)$ being integrated of order 2, i.e., its differences $x(\tau)-x(0)$ would not possess a stationary distribution while its second differences $x(2\tau)-2x(\tau)+x(0)$ would. Thus we take $\kappa\to 0$ while $\lambda$ remains finite. The process $x(t)$ is now characterized by the mean squared displacement function [^19] Alternatively $\langle\dot{x}(0)\dot{x}(\tau)\rangle$, which is one-half the second derivative of the mean squared displacement (Eq. (85)).:
(404) $V(\tau):=\langle(x(\tau)-x(0))^{2}\rangle=2[C(0)-C(\tau)],$
where the last equality holds for a stationary process. We may evaluate $V(\tau)$ using this equality and then taking the limit as $\kappa\to 0$. In this limit, $\mu_{+}=\lambda-\kappa/\lambda$ and $\mu_{-}=\kappa/\lambda$, and hence:
(405) $V(\tau)=\frac{2}{\lambda}\left[{\left({D-\frac{D^{\prime}}{\lambda^{2}}}\right)(1-e^{-\lambda\tau})+\frac{D^{\prime}}{\lambda}\tau}\right].$
For a linear Markov process, $V(\tau)=2D\tau$. We see that $\ddot{V}(0)=0$ if and only if $V(\tau)=2D\tau$. Furthermore, $\ddot{V}(\tau)$ has the same sign as $\ddot{V}(0)$ and thus we may instead test using an integral of the velocity autocovariance function.
Of course, the aforementioned tests can only detect the particular form of non-Markovianity occurring for the simple model investigated, and not more complicated models. However, the results presented indicate that these are reasonable tests for detecting non-Markovianity. This approach can be used not only for the “raw” state variables, but also for the quantities characterizing a Markov process, i.e., quadratic functions, angular momenta, and diffusivities.
@@ 10. Conclusions
We have used Koopman eigenfunctions to facilitate calculations in the case of integrated variables (written in terms of martingales) and nonlinear drift. In the latter case, we might consider these to be new coordinates under a nonlinear coordinate transformation, as these quantities obey a law of linear drift. However, the time-reversed dynamics of Koopman eigenfunctions do not necessarily obey a law of linear drift [^20] The time-reversed dynamics has the same eigenvalues of the Perron–Frobenius and Koopman operators. To see this, start from the Kolmogorov backward equation [22]: $\partial p(\mathbf{x},t\mid\mathbf{x}^{\prime},t^{\prime})/\partial t^{\prime}=-\mathcal{K}p(\mathbf{x},t\mid\mathbf{x}^{\prime},t^{\prime})$ for $t^{\prime}<t$, where $\mathcal{K}$ acts on $\mathbf{x}^{\prime}$. Now use Bayes’s rule to write $p(\mathbf{x}^{\prime},t^{\prime}\mid\mathbf{x},t)=p(\mathbf{x},t\mid\mathbf{x}^{\prime},t^{\prime})p(\mathbf{x}^{\prime})/p(\mathbf{x})$ and therefore $\partial p(\mathbf{x}^{\prime},t^{\prime}\mid\mathbf{x},t)/\partial t^{\prime}=-\mathcal{L}p(\mathbf{x}^{\prime},t^{\prime}\mid\mathbf{x},t)$ ($\mathcal{L}$ acting on $\mathbf{x}^{\prime}$) where $(\mathcal{L}f)(\mathbf{x}):=p(\mathbf{x})\mathcal{K}[f(\mathbf{x})/p(\mathbf{x})]$. It is easily seen that $\mathcal{L}$ is the adjoint of $\mathcal{P}$ with respect to the inner product defined by $(f,g):=\int\mathrm{d}\mathbf{x}\,f(\mathbf{x})^{*}g(\mathbf{x})/p(\mathbf{x})$, and therefore $\mathcal{L}$ has the same eigenvalues as $\mathcal{P}$. (The time-reversed process is also Markov, as the backward time-evolution of the probability density depends only on its current values.). (For example, in the two-dimensional stationary dynamics presented in section 5, if $x$ and $y$ are Koopman eigenfunctions, then third-order covariance functions would be time-symmetric, which is not necessarily the case.) Hence, our use of Koopman eigenfunctions is restricted to being a calculational tool.
We have introduced criteria for judging quantitative significance, both for interesting effects such as non-Gaussianity or time-irreversibility, and for comparing experiment with theory. Our analysis takes linear Gaussian models as a reference point and utilizes second- and third-order covariance functions. Due to moment closure problem, our analysis is limited to asymptotic expansion in the drift nonlinearity. For strongly nonlinear drift, other methods may be more suitable [24, 2].
@@
We expect that the framework presented in this work will be useful in analyses of high-dimensional stochastic systems. We note that extensions to our analysis are possible only up to fourth order (cubic polynomial for drift and quadratic polynomial for diffusion). Afterwards, the analysis procedure may fail (see Appendix B). Besides, analysis of moments beyond fourth order may be infeasible in biological applications.